1x + 0.99y + 0.13z =0.5(x+y+z)
0.99x + 0.05y + 0.02z = 0.24 (x+yxz)

Find
X : Y : z....??

Answer by Alan3354(69443)

1x + 0.99y + 0.13z =0.5(x+y+z)
0.99x + 0.05y + 0.02z = 0.24 (x+yxz)
-----------
Confirm your entry.
0.24 (x+yxz) doesn't seem right.